OpenStudy (anonymous):

Help please ! write a real world situation that could be represented by the system y=3x+10 and y=5x+20

OpenStudy (anonymous):

it gives 5x+20=3x+10 witch means 5x-3x=10-20 means 2x=-10 so x=-5